Compute the 2018 CT tax liability using the info below. Assume married filing jointly. Age 66

W-2 income $85,000
Teachers pension( teachers retirement board)$15,000
social security $30,000
military pension $22,000
INTEREST:
Bank $2,000
state of CT Municipal Bond $3,000
State of Massachusetts Municipal Bond $1,000
Common wealth of Puerto Rico tax exempt $5,000
U.S savings Bond $4,000
Loss on the sale of CT Municipal Bond ($2,000)
Schedule C income (construction) $30,000
ADJUSTED GROSS INCOME $186,000

OTHER INFORMATION:
R.E. taxes paid on home: $3500
CT withholding taxes deducted from payroll $4,000
personal property tax on vehicles $500
Chet contribution for college $12,000.
Answer #1

Income = 85,000

Teachers's pension = 15,000 is exempted from taxation. (age above 65 tax exemption is 15000, when file jointly it is 30,000)

Social security = $30,000

Military pension = $22,000 is fully exempted from taxation

Bank interest = 2000

State of CT Municipal Bond = $3,000 (Exempted from taxation)

State of Massachusetts Municipal Bond = $1,000 (exempted from taxation)

U.S savings Bond = $4,000

Schedule C income (construction) = $30,000

Adjusted gross income = $186,000

Net Income = Income + Social security + Bank interest + U.S. saving bond + Schedule C income + Adjusted gross income

= 85,000 + 30,000 + 2,000 + 4,000 + 30,000 + 186,000 = 337,000

Expenses = R.E. taxes paid on home + Tax withholding + Personal property tax on vehicles + Chet contribution for college

= 3500 + 4,000 + 500 + 12,000 = 20,000

Taxable income = 337,000 - 20,000 = 317,000

Tax liability = 6.90% of 317,000 = 21,873

Common wealth tax exempt = $5,000

So net tax liability is = 21,873 - 5000 =16,873